OpenStudy (anonymous):

what percent of 50 is 18?

OpenStudy (katrinakaif):

What percent = n Of = mulitply 50 is converts to being = 18 n x 50 = 18

OpenStudy (anonymous):

36%

OpenStudy (anonymous):

36% 18/50

OpenStudy (anonymous):

18/50 = 0.36 == 36%

OpenStudy (katrinakaif):

Correct =)

hero (hero):

\[\frac{is}{of}=\frac{\%}{100} \]

OpenStudy (anonymous):

whatsteps could i use to get my answer?

OpenStudy (anonymous):

18/50 = X/100 THEN SOLVE FOR X

OpenStudy (anonymous):

CROSS MULTIPLY

OpenStudy (katrinakaif):

What percent of 50 is 18? ---> n x 50 = 18 n = 18/50 n = .36 Convert this into a percentage because the question asks "WHAT PERCENTAGE..." Move the decimals two places to the left .36 ---> 36%

OpenStudy (anonymous):

just divide the smaller one by the bigger one then multiply by 100. say you get 73/ 100 on a test this equals 73 %. (73/100*100)

OpenStudy (anonymous):

thank u you guys

OpenStudy (katrinakaif):

Anytime =)